# Proficient Graph Neural Network Design by Accumulating Knowledge on Large Language Models

### v2

@@ B.2.2 Model Refinement and Short-run Efficiency
Appendix C Ablation Studies
To delve deeper into the pivotal designs of DesiGNN, we conducted ablation and hyperparameter studies on three key modules, including Graph Understanding (Section 3.1), Knowledge Retrieval (Section 3.2), and GNN Model Suggestion and Refinement (Section 3.3).
C.1 Main Ablation Study
@@
We present the main ablation study on the Knowledge Retrieval and the GNN Model Suggestion and Refinement module in Table 11, quantifying the impact of the retrieved knowledge and each mechanism in our framework. The results from the initial model suggestion stage underscore the importance of designing a proper Graph Understanding method, demonstrating that relying solely on filtered graph properties in dataset descriptions outperforms any use of descriptional inputs. In the short-run performance evaluation after 30 model validations, the Re-rank mechanism significantly improves the knowledge-driven model refinement process when the original order of the Knowledge Pool $\mathcal{K}$ is incorrect. Additionally, the Model Promotion mechanism, which simulates the strategy of human experts refining a model based on accumulated knowledge, plays a crucial role in enhancing the efficacy of model refinement. Lastly, the Directional Exploration mechanism is empirically beneficial, as it leverages the most relevant model design knowledge and the in-context learning ability of LLMs to further refine the best candidate model promoted.
Table 12: Average ranks of different combinations across datasets. A higher rank (with 1 being the highest) corresponds to better performance and lower variability in the initial model proposals.
Combined Rank
Top-$N_{s}$ Desc. g=0 g=2 g=4 g=6 g=8 g=10 g=16
$N_{s}$=1 True 2.250 2.312 2.250 2.125 2.437 2.625 2.312
False 2.562 2.125 2.562 1.812 2.187 2.250 2.125
$N_{s}$=2 True 2.500 2.187 2.625 1.937 1.687 2.312 2.750
False 2.625 2.375 2.812 2.000 1.500 2.187 2.250
$N_{s}$=3 True 2.500 1.937 2.875 1.875 1.937 2.187 2.000
False 2.125 1.687 2.000 1.750 1.500 1.3750 1.687
C.2 Ablation Studies on Graph Understanding
In this section, we study the combined effect of the number of selected properties $N_{f}$ and the usage of descriptional inputs in understanding and comparing graph datasets, under varying sizes $N_{s}$ of the Knowledge Pool, which influences the number of initial proposals suggested by LLMs (Section 3.3). The average rankings of different combinations across datasets in Table 12 illustrate that while the optimal $N_{f}$ varies with different $N_{s}$ values, integrating descriptional inputs often diminishes performance, particularly when $N_{f}>6$ and $N_{s}=3$. This finding supports our notion of “artificial hallucinations,” where descriptive inputs can hinder the understanding of tasks when sufficient graph properties are provided. From another aspect, this result also validates our idea of analyzing graph properties to design tailored GNN models, which came from insights documented in the literature.
Table 13: The distribution of the best initial model among the Top-3 benchmark.
Is Best Initial Model (%)
Scenario Best Second Third
Descriptional Only 30.0 50.0 20.0
$N_{f}=6$ 50.0 22.5 27.5
$N_{f}=8$ 52.5 37.5 10.0
$N_{f}=10$ 42.5 40.0 17.5
Table 14: Average performance rank of transferring the Top-$N_{m}$ model designs.
Average Rank
Top-$N_{m}$ $N_{m}$=1 $N_{m}$=10 $N_{m}$=20 $N_{m}$=30 $N_{m}$=40 $N_{m}$=50
2.7077 2.8750 2.8438 2.6875 2.9231 2.8769
C.3 Ablation Studies on GNN Model Suggestion
This study extends from Section 3.1 and quantifies the advantage of using the two $N_{s}=2$ or three $N_{s}=3$ most similar benchmarks instead of solely the most similar one $s=1$ (no tolerance). Table 13 shows the distribution of the best initial proposal stems from the first, second, or third most similar benchmark. For $N_{f}=8$ without descriptive inputs, approximately $53\%$ of trials identify the optimal initial proposal from the most similar benchmark, with only $10\%$ benefiting from the third. This data supports a balanced trade-off between performance and efficiency. Notably, this study also serves as the motivation for us to design the Re-rank Mechanism in Model Proposal Refinement (Appendix A.4).
Moreover, we explore the optimal number $N_{m}$ of top-performing examples from each benchmark when establishing the Knowledge Pool and assess the impact of using poor examples. Table 14 shows that $N_{m}=30$ yields the best average performance across all dataset combinations ($2160$ records), with further analysis indicating that poor examples do not significantly affect performance outcomes (T-statistic $=-0.1262$, P-value $=0.8998$). We thus adopt $N_{m}=30$ without bad examples for an appropriate token length of the LLM prompt.
C.4 Ablation Studies on GNN Model Refinement
We investigate the impact of using the $\mathcal{K}_{1}$ Knowledge Pool as context for LLMs and the Re-rank Mechanism that is based on the performance ranking of initial proposals. We also study the best number of candidates and the effect of adopting different $N_{s}$ numbers of knowledge bases to guide the controlled search space exploration. Table 15 displays T-statistics for various configurations of Knowledge Pool usage and Re-rank Mechanism. For the row of w/o candidate, the True/False columns stand for w/o re-rank, and vice versa. While statistical significance is not reached (p-value $>0.05$), certain trends emerge:
  • When $N_{s}=3$ knowledge bases are used for a broader perspective, employing a Re-rank Mechanism consistently adds value. This pattern is intuitive as the Re-rank Mechanism will become more influential as the number of knowledge bases increases.
  • Broader exploration with $N_{s}=3$ and a larger number of candidates (e.g., $30$) is likely effective, as we also have a directional exploitation mechanism to refine the diversity of potentially immature solutions.
Given the complexity of these patterns, we further assess the performance ranking of GNN Model Refinement hyperparameter combinations across benchmarks to determine the most effective settings. Table 16 suggests that the best hyperparameter combination is $N_{s}=3$ and $N_{m}=30$ with both Re-rank Mechanism and knowledge-driven refinement strategy enabled.
Table 15: T-statistic of the ablation on the usage of knowledge-driven Exploration and Re-rank Mechanism with different candidate number and $N_{s}$.
Candidate Number= 10 20 30
w/o Re-rank or w/o Exploration
Ablation $N_{s}$ True False True False True False
w/o Exploration 2 -0.280 0.582 0.512 1.977 0.557 0.566
3 0.863 0.072 0.055 -0.164 -0.054 -0.358
w/o Re-rank 2 0.647 1.383 -0.334 1.232 -2.092 -0.852
3 -0.116 -0.808 -0.593 -0.591 -0.751 -1.904
Table 16: Average performance rank of the GNN Model Refinement hyperparameter combinations across benchmark.
Candidate Number= 10 20 30
w/o Re-rank
$N_{s}$ w/o Exploration True False True False True False
2 True 7.33 12 7.83 9 13.5 9.17
False 7.16 14.5 11.7 17.33 16.17 11.83
3 True 11.7 11.17 11.83 9.5 15.33 7.33
False 12.33 10.17 11.17 10.17 12.83 7
Appendix D Case Studies
D.1 Case Studies: Lack of Prior Knowledge in LLMs
The first study verifies the inherent knowledge gap illustrated in Figure 1(a): lack of prior knowledge in LLMs about the top-performing models of benchmark datasets within the NAS-Bench-Graph [35] model space (Section A.3). The first two columns in Table 20 are cases where the descriptive inputs or graph properties are directly sent to LLMs for model suggestions. The results demonstrate that LLMs are only able to suggest commonplace layer connections and operations based on user-provided descriptional inputs (left panel) or key graph properties (middle panel). In fact, we have tested all the benchmark datasets in NAS-Bench-Graph [35] with descriptive inputs, and we found only two macro lists and three operation lists that LLM would recommend regardless of the datasets:
  • Architecture: [0, 1, 2, 3] and [0, 0, 1, 3].
  • Operations: [’gcn’, ’gat’, ’sage’, ’skip’], [’gcn’, ’gat’, ’sage’, ’gcn’], and [’gcn’, ’gat’, ’gin’, ’sage’],
After providing the tailored knowledge retrieved from our framework to LLMs (right panel), it can be clearly seen that the model it recommends is specialized and has obvious performance improvements.
D.2 Case Studies: Artificial Hallucination in LLMs
The second case study in Table 21 and Figure 1(b) examines the phenomenon of “artificial hallucination”, or the external noise issue, in LLMs when comparing the similarities between the unseen dataset and the benchmark datasets, as detailed in Section 3.2. Figure 5 (leftmost) illustrates that the empirically most similar benchmark datasets to PubMed are CS, Physics, and Citeseer, in descending order of similarity. When employing our Graph Understanding method in Section 3.1, which leverages confident graph properties (right panel), the top three most similar datasets are Citeseer, CS, and Physics (align well with empirical ground truth). However, the current practice that relies solely on descriptive inputs (left panel) identifies Cora, Citeseer, and ogbn-arxiv as the top three, which does not align with the empirical ground truth. This discrepancy arises because LLMs overly rely on the shared characteristic of being citation graphs, assuming that citation datasets like PubMed, Cora, Citeseer, and ogbn-arxiv should have similar model design preferences. This case study demonstrates that relying solely on external descriptive inputs is insufficient to capture the similarities between datasets, which can overwhelm other pertinent information, leading to inaccurate task similarity assessment and incorrect knowledge retrieval.
Table 17: Performance comparison of initial model suggestion with different LLMs as meta-controller.
@@ D.2 Case Studies: Artificial Hallucination in LLMs
Table 18: Performance comparison of model refinement with different LLMs as meta-controller.
Dataset Cora Cit. Pub. CS Phy. Pho. Com. arX.
@@
Table 19: Performance comparison of different prompt designs on GPT4GNAS.
Dataset Cora Cit. Pub. CS Phy. Pho. Com. arX.
Original 81.31 69.43 76.90 90.44 92.12 92.21 83.96 71.67
Our Prompt 81.38 70.06 77.03 90.22 92.16 92.08 83.78 71.72
D.3 Case Studies: Different LLMs
@@
To study the effect of different LLMs as the meta-controller, we adopt an open-source LLM, Llama2-13b, to replace the GPT-4 meta-controller we originally employed in our paper. Table 17 shows the comparative results (averaged over five runs) on the initial model proposal, while Table 18 shows the comparative results (averaged over five runs) after 30 rounds of model validations. We can see that replacing GPT-4 with Llama2 results in a slight degradation of the performance of recommended models in both settings. This degradation is more consistent after the refinement of the model design. This suggests that GPT-4 is superior at proposing a reasonable next step from the trajectory and extra knowledge (optimization tasks), consistent with the existing study [53]. Notably, because the initial model proposal phrase relies more on the quality of our elicited knowledge, Llama2 can also achieve close performance to GPT-4.
D.4 Case Studies: Prompt Design Impact
We studied whether fine-tuning in prompt engineering has a positive or negative effect on our work. Since we did not adopt any prompt fine-tuning approach, we added an alternative experiment to compare whether there is any consistent difference in the performance of LLM-based baseline GPT4GNAS [43] between using our prompt design and its original design. This includes adding topological graph descriptions and replacing model space descriptions and general instructions. No knowledge-related design from our method is being transferred because it is our key novelty. After prompt replacement, GHGNAS [12] became the same as GPT4GNAS [43], so we only reported one of them. Table 19 below shows the comparative results (averaged over five runs) after 30 rounds of refinements. We can see that after replacing the prompt design of the baseline with our design, the impact on the quality of the recommended models is inconsistent. This study justifies that our method does not improve performance by fine-tuning the prompt; instead, the main source of our performance improvement arises from our tailored model design knowledge.
Figure 8: Short-run performance trajectories of DesiGNN compared to all the automated baselines after validating 1-30 proposals.
Figure 9: Short-run performance trajectories of DesiGNN compared to all the automated baselines after validating 1-30 proposals.
Table 20: Example LLM Response on Recommending Model for PubMed dataset
Section Response (w/ Descriptional Only) Response (w/ Properties Only) Response (w/ Knowledge)
Decision For the unseen dataset: (Architecture: [0, 0, 1, 3], Operations: [’gcn’, ’gat’, ’sage’, ’gcn’]) For the unseen dataset: (Architecture: [0, 0, 1, 3], Operations: [’gcn’, ’gat’, ’sage’, ’gin’]) For the unseen dataset: (Architecture: [0, 0, 0, 1], Operations: [’gcn’, ’gat’, ’arma’, ’graph’])
Reason 1. Architecture Choice ([0, 0, 1, 3]): …… 2. Operations Choice ([’gcn’, ’gat’, ’sage’, ’gcn’]): - First Operation - GCN (’gcn’): The Graph Convolutional Network (GCN) is effective for aggregating neighborhood information, which is essential in a citation graph where the connectivity pattern (citations) plays a crucial role in defining the node (paper) features. - Second Operation - GAT (’gat’): The Graph Attention Network (GAT) introduces an attention mechanism, allowing the model to focus on the most relevant parts of the input graph structure. This is particularly useful in citation graphs for emphasizing influential papers. - Third Operation - GraphSAGE (’sage’): GraphSAGE is adept at sampling and aggregating features from a node’s neighborhood, which can help in efficiently handling the local neighborhood structures and diversity in node degrees typical in citation graphs. - Fourth Operation - GCN (’gcn’): Another layer of GCN helps to further refine the feature representation by aggregating higher-order neighborhood information, which can be crucial for capturing the broader context in the graph. This combination and sequence of operations are designed to leverage both the local and global structural information in the graph, which is critical for a citation graph where both direct and indirect citations (connections) can influence the classification of a paper into topics. The use of both GCN and GAT helps in effectively capturing different levels of granularity in neighbor interactions, while GraphSAGE allows for efficient computation by focusing on a sampled subset of neighbors. The architecture’s mix of parallel and sequential processing ensures a comprehensive feature transformation pathway, enhancing the model’s ability to learn complex patterns in the data. 1. Architecture Choice ([0, 0, 1, 3]): …… 2. Operations Choice ([’gcn’, ’gat’, ’sage’, ’gin’]): - ’gcn’ (Graph Convolutional Network layer): Effective for capturing neighborhood information through a simple yet powerful convolutional approach. Given the average degree and edge count, ’gcn’ can efficiently aggregate local neighborhood features without being too computationally expensive. - ’gat’ (Graph Attention Network layer): Utilizes attention mechanisms to weigh the importance of nodes’ neighbors. This is particularly useful in graphs where nodes vary significantly in their roles and influence, as indicated by the centrality measures. - ’sage’ (GraphSAGE): Samples and aggregates features from a node’s neighborhood, which can handle the graph’s density and scale effectively. This method is adaptive and can learn from different-sized neighborhoods, which is beneficial given the variability in node connectivity. - ’gin’ (Graph Isomorphism Network layer): Designed to capture more nuanced graph structural information, which can be crucial for understanding deeper patterns in the graph, especially given the non-trivial graph diameter and clustering coefficient. This combination and sequence of operations are designed to first broadly and then selectively refine the feature representation of the nodes, considering both local and global structural properties of the graph. This should provide a robust model capable of handling the complexities and variabilities in the graph metrics provided. 1. Architecture Choice: …… 2. Operation Selection: …… - ’arma’ (Auto-Regressive Moving Average): ARMA layers have been chosen in various top models, suggesting their utility in capturing more complex dependencies in graph data than standard convolutional layers. - ’graph’ (k-GNN): The inclusion of a k-GNN operation in the final step can extend the model’s capability to capture higher-order interactions, which might be crucial for the specific characteristics of the unseen dataset. 3. Combination of Operations: The combination of GCN, GAT, and ARMA before a final GRAPH operation provides a robust sequence of transformations. Starting with GCN allows for a broad capture of neighborhood information, followed by the selective attention mechanism of GAT and the complex filtering of ARMA, culminating in a k-GNN layer that can capture higher-order dependencies. This sequence is designed to progressively refine and enhance the feature representations. 4. General Observations from Top Models: The selected operations and architecture are among the most recurrent in the top-performing models from similar datasets. This suggests that they are likely to be effective in handling the types of graph-structured data encountered in these benchmarks, making them a prudent choice for the unseen dataset. By synthesizing these observations, the recommended model design aims to balance complexity and depth of feature extraction and transformation, which is crucial for effectively processing graph-structured data. This design is expected to perform well on the unseen dataset by leveraging proven elements from similar successful models.
Perf. 75.40 71.56 76.60
Table 21: Example LLM Response on Dataset Comparison for PubMed dataset
Benchmark w/ Descriptional Inputs w/ Confident Graph Properties
Cora Similarity Score: 0.8 Reason: Both the unseen dataset and Cora are citation graphs with papers as vertices and citations as links. Both use bag-of-words features and have ground-truth topics as labels. The main differences are in the number of features and classes, with Cora having more features and classes. Similarity Score: 0.1 Reason: The unseen dataset has a significantly higher graph diameter, average closeness centrality, and average betweenness centrality compared to Cora. Additionally, the general graph metrics such as edge count, average degree, and density are quite different, indicating a low similarity.
Citeseer Similarity Score: 0.75 Reason: Similar to Cora, Citeseer is also a citation graph with similar characteristics to the unseen dataset. The differences are more pronounced in the number of features and classes, with Citeseer having significantly more features. Similarity Score: 0.6 Reason: Citeseer has a closer match in terms of graph diameter and average closeness centrality with the unseen dataset. The average betweenness centrality and general graph metrics like edge count and average degree also show some similarities, suggesting a moderate level of task similarity.
CS Similarity Score: 0.2 Reason: CS is a co-authorship graph, differing significantly in the type of vertices (authors vs. papers) and the nature of features (keywords vs. bag-of-words). The task and evaluation metric are similar, but the structural differences lower the similarity score. Similarity Score: 0.3 Reason: CS shares a similar graph diameter with the unseen dataset but differs significantly in average closeness centrality, average betweenness centrality, and general graph metrics such as edge count and average degree, indicating a moderate to low similarity.
Physics Similarity Score: 0.2 Reason: Similar to CS, Physics is a co-authorship graph with authors as vertices and co-author relationships as links, which is structurally different from the unseen dataset’s citation graph nature. Similarity Score: 0.2 Reason: Physics has a similar graph diameter but differs significantly in average closeness centrality, average betweenness centrality, and other general graph metrics like edge count and average degree, suggesting a low similarity.
Photo Similarity Score: 0.3 Reason: Photo is a co-purchase graph, which is different in nature from a citation graph. However, both use bag-of-words features and have a similar number of features, slightly increasing the similarity score. Similarity Score: 0.1 Reason: Photo, while having a similar graph diameter, differs significantly in average closeness centrality, average betweenness centrality, and general graph metrics such as edge count and average degree, indicating a low similarity.
Computers Similarity Score: 0.3 Reason: Similar to Photo, Computers is also a co-purchase graph. The similarity is slightly higher due to a closer number of features and classes compared to the unseen dataset. Similarity Score: 0.1 Reason: Computers has a similar graph diameter but differs significantly in average closeness centrality, average betweenness centrality, and general graph metrics like edge count and average degree, indicating a low similarity.
ogbn-arxiv Similarity Score: 0.7 Reason: ogbn-arxiv is a citation graph similar to the unseen dataset, but it uses word embeddings for features instead of bag-of-words, and has a significantly different number of classes. Similarity Score: 0.2 Reason: ogbn-arxiv shares a similar graph diameter but differs significantly in average closeness centrality, average betweenness centrality, and general graph metrics such as edge count and average degree, suggesting a low similarity.
ogbn-proteins Similarity Score: -0.5 Reason: ogbn-proteins is structurally and fundamentally different from the unseen dataset, being a protein association graph with a multi-label binary classification task and ROC-AUC as the evaluation metric. Similarity Score: 0.1 Reason: ogbn-proteins has a different graph diameter and significantly different average closeness centrality, average betweenness centrality, and general graph metrics such as edge count and average degree, indicating a very low similarity.

### v3

@@ 4.2. Main Results
Figure 6. Task similarities computed by empirical performance (target), two meta-based baselines, and ours. Kendall’s $\tau$ correlation coefficient with the empirical ground truth is shown in parentheses.
Table 4. Hit rate $\delta^{*}$ across unseen datasets.
Top-s Benchs $N_{s}$=1 $N_{s}$=2 $N_{s}$=3
Kendall 0 0.25 0.625
Overlap 0 0.375 0.625
Property Similar. (w/o LLMs) 0.375 0.5 0.625
Description Only (with LLMs) 0.262 0.5 0.575
Property Only (with LLMs) 0.375 0.6 0.725
Both (with LLMs) 0.275 0.5 0.6
Model Refinement and Short-run Efficiency. Building on the initial proposals, DesiGNN can rapidly refine models to outperform all manually designed and automated GNNs across 11 datasets, demonstrating its efficient use of meta-knowledge to exploit the high potential range. Figure 5 highlights DesiGNN’s short-run efficiency, achieving significant performance improvements with fewer model proposal validations than automated baselines. Unlike methods that require extensive validation to identify feasible models, DesiGNN emulates human-like refinement strategies, reducing computational costs and accelerating optimization. Specifically, Table 2 shows that other automated methods require significantly more model validations to match DesiGNN’s performance after just 10 validations. These results confirm that DesiGNN effectively addresses computational inefficiencies by leveraging accumulated knowledge to enable immediate feedback and rapid optimization, delivering consistently superior performance under limited budgets.
4.3. Effective Knowledge via Task Similarity
To assess whether our unsupervised task similarity in Equation 5 reflects the probability of successful knowledge transfer defined in Definition 3.1, we quantify the hit rate, defined as the tighter probability $\delta^{*}$ that the empirically most relevant benchmarks (i.e., those maximizing $\mathcal{H}(\{\theta^{*}_{im}\},\cdot,G^{u})$) are included in the Top-$N_{s}$ benchmarks retrieved by $\mathcal{S}$ ($N_{s}$ is the tolerance). A higher hit rate directly supports the probabilistic claim in Definition 3.1, where $\mathcal{S}$ correlates with the likelihood of transferring high-performing models from $G^{i}$ to $G^{u}$. As shown in Table 4, our unsupervised task similarity based on graph properties achieves the highest hit rate across all tolerance settings, outperforming methods that rely on Kendall or Top 5% similarities, as well as purely statistical property similarity (without LLMs to conduct adaptive knowledge elicitation and alignment). This demonstrates that our performance-oriented $\mathcal{S}$, derived from key graph properties and adaptive weighting by LLMs, effectively captures the meta-knowledge required for data-aware model design. Besides, we observed a negative effect when incorporating descriptive inputs into similarity assessment, suggesting that some semantic properties (e.g., being a citation graph) are less relevant to designing the model architecture (see the case study in Figure 1(b)). We further visualize the pairwise similarities between graphs in Figures 3(b) and  6, showing that our unsupervised task similarity closely aligns with empirical performance similarity (the highest Kendall’s $\tau$ correlation). This further validates the claim that $\mathcal{S}$ approximates a pseudo-likelihood $\mathbb{P}(\mathcal{F}(G^{u})|\theta_{i}^{*})$ well, reinforcing its role in supporting the Bayesian-inspired framework in Section 3.2.
4.4. Main Ablation Study
@@
We present the main ablation study on the Knowledge Retrieval and the GNN Model Suggestion and Refinement module in Table 3. The results from the initial model suggestion stage underscore the importance of designing an effective Graph Understanding method, demonstrating that relying solely on filtered graph properties in dataset descriptions outperforms using descriptional inputs. In the short run, after 30 model validations, the Re-rank mechanism significantly improves the knowledge-driven model refinement process when the original order of the Knowledge Pool $\mathcal{K}$ is incorrect. Additionally, the Model Promotion mechanism, which simulates the strategy of human experts refining a model based on accumulated knowledge, plays a crucial role in enhancing the efficacy of model refinement. Lastly, the Directional Exploration mechanism is empirically beneficial, as it leverages the most relevant model design knowledge and the in-context learning ability of LLMs to further refine the best candidate model promoted.
Table 5. Different LLMs for model suggestion and refinement.
@@
Model Refinement
@@
4.5. Case Studies
@@
Different LLMs. To study the effect of different LLMs as meta-controllers, we use an open-source Llama2-13b to replace the GPT-4 meta-controller. Table 5 shows the comparative results on the initial model proposal and refinement. We observed that replacing GPT-4 with Llama2 results in a slight degradation of the performance of recommended models in both settings, and is more consistent in model refinement. This suggests that GPT-4 is superior at proposing a reasonable next step from the trajectory and extra knowledge, consistent with the existing study (Yang et al., 2023). Notably, because the initial model proposal phrase relies more on the quality of our elicited knowledge, Llama2 can also achieve close performance to GPT-4.
Lack of Prior Knowledge in LLMs. The inherent knowledge gap illustrated in Figure 1(a) refers to the lack of prior knowledge in LLMs about the top-performing models of benchmark datasets within the NAS-Bench-Graph (Qin et al., 2022) model space. We analyze LLM responses in which the descriptive inputs or graph properties were directly sent to the LLMs for model suggestions. The results demonstrate that LLMs are only able to suggest commonplace layer connections and operations based on user-provided descriptive inputs or key graph properties. We have tested all benchmark datasets in NAS-Bench-Graph and found that only two macro lists and three operation lists would be recommended by LLMs regardless of the datasets: (1) Architecture: [0, 1, 2, 3] and [0, 0, 1, 3]; (2) Operations: [’gcn’, ’gat’, ’sage’, ’skip’], [’gcn’, ’gat’, ’sage’, ’gcn’], and [’gcn’, ’gat’, ’gin’, ’sage’]. After providing tailored knowledge derived from our framework, the LLM-recommended model becomes specialized and shows significant performance improvements.
Artificial Hallucination in LLMs. In Figure 1(b), we examine the phenomenon of “artificial hallucination,” or the external noise issue, in LLMs when comparing the similarities between unseen and benchmark datasets (Section 3.2). Figure 6 (leftmost) illustrates that the empirically most similar benchmark datasets to PubMed are CS, Physics, and Citeseer, in descending order of similarity. When employing our Graph Understanding method in Section 3.1, which leverages confident graph properties, the top three most similar datasets are Citeseer, CS, and Physics (align with the empirical ground truth). However, the current practice that relies solely on descriptive inputs identifies Cora, Citeseer, and ogbn-arxiv as the top three, which does not align with the empirical ground truth. This discrepancy arises because LLMs overly rely on the shared characteristic of being citation graphs, assuming that citation datasets like PubMed, Cora, Citeseer, and ogbn-arxiv should have similar model design preferences. This reveals that relying solely on external descriptive inputs is insufficient to capture the similarities between datasets, which can overwhelm other pertinent information, leading to inaccurate task association and knowledge retrieval.
5. Conclusion
We present DesiGNN, a knowledge-centered framework that demonstrates deep expertise in proficiently designing data-aware GNNs and eliminates external noise by systematically converting accumulated model design experience into fine-grained meta-knowledge well suited to meta-learning with LLMs. Leveraging a performance-oriented task similarity grounded in empirically validated graph properties and adaptive knowledge elicitation via LLMs, DesiGNN bridges the gap between literature insights and practical, data-specific model deployment. DesiGNN integrates graph understanding, effective knowledge retrieval, and strategic model suggestion and refinement under a Bayesian perspective, enabling tailored GNN designs for diverse graphs in a single shot or with a few attempts. Extensive experiments demonstrate that DesiGNN expedites the design process by delivering top-tier GNN designs within seconds and achieving superior performance at minimal cost.
Acknowledgements.
Lei Chen’s work is partially supported by National Key Research and Development Program of China Grant No. 2023YFF0725100, National Science Foundation of China (NSFC) under Grant No. U22B2060, Guangdong-Hong Kong Technology Innovation Joint Funding Scheme Project No. 2024A0505040012, the Hong Kong RGC GRF Project 16213620, RIF Project R6020-19, AOE Project AoE/E-603/18, Theme-based project TRS T41-603/20R, CRF Project C2004-21G, Key Areas Special Project of Guangdong Provincial Universities 2024ZDZX1006, Guangdong Province Science and Technology Plan Project 2023A0505030011, Guangzhou municipality big data intelligence key lab, 2023A03J0012, Hong Kong ITC ITF grants MHX/078/21 and PRP/004/22FX, Zhujiang scholar program 2021JC02X170, Microsoft Research Asia Collaborative Research Grant, HKUST-Webank joint research lab, 2023 HKUST Shenzhen-Hong Kong Collaborative Innovation Institute Green Sustainability Special Fund, from Shui On Xintiandi and the InnoSpace GBA, and HKUST(GZ) - CMCC(Guangzhou Branch) Metaverse Joint Innovation Lab under Grant No. P00659. Jiachuan Wang’s work is supported in part by JST CREST (JPMJCR22M2). Shimin Di’s work is supported by National Science Foundation of China (NSFC) under Grant No. 62506075.